Len Mast is married with three allowances. His weekly gross pay is $1,250. He will not reach the maximum taxable salary for Social Security purposes.

1. What is the amount that he will have deducted for Social Security taxes?

2. What is the amount that he will have deducted for Medicare taxes?

3. Using the percentage method, what is the amount that he will have deducted for FIT (Federal Income Tax)?

SOCIAL SECURITY TAX RATE: 6.2% MEDICARE TAX RATE: 1.45% MAXIMUM TAXABLE SALARY: $110,100 Payroll Period Weekly Biweekly SemimTABLE 2—BIWEEKLY Payroll Period (a) SINGLE person (including head of household) (b) MARRIED person, If the amount of wages (a 4. What is his net pay?
Answer #1
1
Deduction for Social Security taxes 77.50 =1250*6.2%
2
Deduction for Medicare taxes 18.13 =1250*1.45%
3
Gross pay 1250.00
Less: Withholding allowance 219.24 =73.08*3
Taxable pay 1030.76
Deduction for FIT (Federal Income Tax) 114.51 =33.40+(1030.76-490)*15%
4
Gross pay 1250.00
Less: Deductions
Social Security taxes 77.50
Medicare taxes 18.13
Federal Income Tax 114.51
Total Deductions 210.14
Net pay 1039.86